2.   A carbonate buffer of pH 10 is needed by a chemist. How many grams of Na2CO3 is needed to dissolve in 1.5L of 0.2M NaHCO3 to make the buffer? Ka of HCO3- = 4.7 x 10-11



1
Expert's answer
2022-05-12T03:52:05-0400

Given:

pH = 10

[NaHCO3] = 0.2 M

Liters of solution = 1.5 L

Ka of HCO3 = 4.7×10−11


Solution:

The Henderson-Hasselbalch equation is commonly used to calculate the pH of a buffer solution from the concentration of the buffer components:



where pKa is the acid dissociation constant and [base] and [acid] are the base and acid concentrations, respectively.

In this specific reaction, the base is Na2CO3 and the acid is NaHCO3.

Therefore,

10 = −log(4.7×10−11) + log([Na2CO3] / 0.2)

−0.3279 = log([Na2CO3] / 0.2)

0.47 = [Na2CO3] / 0.2

[Na2CO3] = 0.094

Thus, the molarity of Na2CO3 is 0.094 M


Molarity = Moles of solute / Liters of solution

Therefore,

Moles of Na2CO3 = Molarity of Na2CO3 × Liters of solution = (0.094 M) × (1.5 L) = 0.141 mol


Moles = Mass / Molar mass

The molar mass of Na2CO3 is 106 g/mol

Therefore,

Mass of Na2CO3 = (0.141 mol Na2CO3) × (106 g Na2CO3 / 1 mol Na2CO3) = 14.946 g Na2CO3

Mass of Na2CO3 = 14.95 g


Answer: 14.95 grams of Na2CO3 is needed to dissolve